If F(s)=s^{2}+2/s(2s^{2}-7s+5); then lim t→∞f(t)


1
Expert's answer
2021-08-09T08:04:25-0400

"F(s) = \dfrac{s²+2}{s(2s²-7s+5)}"


"= \dfrac{s²+3}{2s³-7s²+5s}"


Using quotient rule

"u = s²+2;\ \ du = 2s\\

v = 2s³-7s²+5s; \ \ dv= 6s²-14s+5"


"F(s) =\dfrac{Vdu-Udv}{V²}"


"F'(s) = \dfrac{2s(2s³-7s²+5s)}{(s²+2)(6s²-14s+5)}"


"=\dfrac{4s⁴-15s³+10s²}{6s⁴-14s³+5s²+12s²-28s-10}"


"=\dfrac{4s⁴-15s³+10s²}{6s⁴-14s³+17s²-28s-10}"